OpenStudy (anonymous):

If a shirt costs $64 after a 20% discount, what was its original price?

OpenStudy (kropot72):

$64 is 80% of the original price. Therefore the original price was:\[\frac{64}{80}\times 100=80\]dollars

OpenStudy (anonymous):

yes thnk u..

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

Let original price =y=100%. Now, given that 80%= $64, it follows that 100%=64/80*100=80 dollars

OpenStudy (anonymous):

yeah